Perpetrators of a crime who have been given the maximum sentence of 10 years are

believed bya fiscal to have been sentenced to no more than 7 years on the average.

random sample of 20of these crime offenders reveals a mean sentence of 7.6 years with a

standard deviation of 2.4 years. Is the fiscal correct in his suspicions at the 0.05 level of

significance?


Step:

1. State the null and alternative hypothesis

concerning the population mean, "\mu" and the type of test to be used.

2.Specify the level of significance "\alpha"

3.State the decision rule

4. Collect the data and perform calculations.

5. Make a statistical decision.

6.State the conclusion


1
Expert's answer
2022-05-10T23:01:19-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu\le7"

"H_a:\mu>7"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

2. Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=19" degrees of freedom, and the critical value for a right-tailed test is "t_c = 1.729133."

3. The rejection region for this right-tailed test is "R = \{t: t> 1.729133\}."

4. The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{7.6-7}{2.4/\sqrt{20}}\approx1.118"


5. Since it is observed that "t=1.118< 1.729133=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, "df=19" degrees of freedom, "t=1.118" is "p=0.138752," and since "p=0.138752>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

6. Therefore, there is not enough evidence to claim that the population mean "\mu" is greater than 7, at the "\alpha = 0.05" significance level.